OpenStudy (risingtotop):

Part A: Solve F = alt= C + 32 for C. (4 points) Part B: Determine the value of C when F = 10 °F. (2 points) Part C: Solve -np - 70 > 40 for n. Show your work. (4 points)

OpenStudy (risingtotop):

@Will.H

OpenStudy (will.h):

Can u write the equation again. It's unclear

OpenStudy (risingtotop):

sorry about that

OpenStudy (will.h):

No worries Alright so solving for C We have the equation F= (9/5)c + 32 At first we have to subtract 32 from both sides F- 32 =(9/5) c Now 9/5 is equivalent to 1.8 that would help to avoid confusion with those fractions so the new shape is F - 32 = 1.8c since 1.8 multiplied then just divide it by both sides (F - 32)/1.8 = C

OpenStudy (will.h):

Make sense mate?

OpenStudy (risingtotop):

hold on im reading it haha

OpenStudy (will.h):

Lol øk

OpenStudy (risingtotop):

ok makes sense

OpenStudy (risingtotop):

are you from britian by any chancec

OpenStudy (will.h):

Yeah. How'd u know

OpenStudy (risingtotop):

Im stalker 101

OpenStudy (risingtotop):

jk because you say mate alot

OpenStudy (will.h):

Lmao

OpenStudy (will.h):

Ohh right. You guys say dude

OpenStudy (risingtotop):

only british people say mate and austrialian

OpenStudy (risingtotop):

yup dude lol

OpenStudy (risingtotop):

alright i understand the lecture you typed :P

OpenStudy (will.h):

Lool Great mate Let's move to B We already have the equation to solve for C C= (F-32)/1.8 Solve for C When F = 10

OpenStudy (will.h):

Solve and I will tell u if u are right

OpenStudy (risingtotop):

lier

OpenStudy (risingtotop):

how old are you

OpenStudy (will.h):

Lol

OpenStudy (will.h):

18

OpenStudy (risingtotop):

O_O

OpenStudy (will.h):

Astonished face?

OpenStudy (risingtotop):

exactly ...

OpenStudy (will.h):

Why lol

OpenStudy (risingtotop):

-12.22222222222222

OpenStudy (risingtotop):

what is an 18 year old doing on openstudy lmao when he could be like idk at a party haha

OpenStudy (risingtotop):

i heard british people have LOTS of parties .....

OpenStudy (risingtotop):

-12.22222222222222 thats what i got

OpenStudy (will.h):

Well am odd And your answer is correct

OpenStudy (risingtotop):

how are you odd?

OpenStudy (will.h):

I don't hang out quite often.

OpenStudy (risingtotop):

yay another me just older hahaha

OpenStudy (risingtotop):

were you stupid like me when you were young btw im 15

OpenStudy (risingtotop):

not in any offensive way

OpenStudy (will.h):

Loool No well I always have been smart that's why I don't do partying quite often

OpenStudy (risingtotop):

i dont and wouldnt i was just never into parties except for for becoming an entrepreneur and living life to the max

OpenStudy (risingtotop):

but thats in another 15 years lmao.... okay back to problem

OpenStudy (will.h):

Lool. Okay back to C

OpenStudy (will.h):

-Np - 70 > 40 Solve for n

OpenStudy (risingtotop):

hoq

OpenStudy (risingtotop):

how

OpenStudy (will.h):

-Np - 70 >40 First add 70 to both sides -Np> 110 Now divide both sides by p -n >110/P We still have to get rid of the negative. So divide by -1 when dividing by negative we reverse the sign n < -110/P Make sense mate?

OpenStudy (risingtotop):

o so same process as if there was a number instead of N

OpenStudy (will.h):

Yeah just normal solve but -n is same as -1n that's all

OpenStudy (risingtotop):

okay i get it

OpenStudy (risingtotop):

thanks

OpenStudy (will.h):

No problem.

OpenStudy (will.h):

Mate

OpenStudy (will.h):

Lol